A vessel of 120 ml capacity current contains a certain mass of gas at 20 degree Celsius and 750 mm pressure the gas was transferred to a vessel whose volume is 180 ml then the pressure of gas is at  20 degree celsius is 

Grade:11

2 Answers

Arun
5 years ago
 = 750 * 120 / 180 = 500mm Hg2 , P2V2=P1V1 =180 ml,  Using boyle's law: , P2=120 ml,  V1=750 mm Hg, V1Dear studentP
Sweety Sharma
5 years ago
Use P2=P1×V1/V2
      P2=750×120/180
      P2=500 mm of Hg
          OR
P1=750/ 760 ; V1=120/1000 litre &  V2=180/1000 litre then  P2=750×120/760×180
                P2=0.657 atm